OpenStudy (anonymous):

9 children can complete a piece of work in 360 days. 18 men can complete the same piece of work in 72 days and 12 women can complete the piece of work in 162 days. In how many days can 4 men, 12 women and 10 children together complete the piece of work?

OpenStudy (anonymous):

you would each to find each persons work. so 360/9=40, each kid is worth 40. 72/18=4, each man is worth 4 days 162/12=13.5, so each women is 13.5

OpenStudy (anonymous):

still i didnt get the nswer

OpenStudy (anonymous):

you would add them together, you need to know 4 men, 12 women, and 10 children together. 1 man is 4 so for men is 16 1 child is 40 so 10 children is 400 1 women is 13.5 so 12 women would be 162...you would just add those together.

OpenStudy (anonymous):

578 ?

OpenStudy (anonymous):

yupp! :D

OpenStudy (anonymous):

578/26 =22. 45

OpenStudy (anonymous):

i hope that made sense!

OpenStudy (anonymous):

why would you divide by 26?

OpenStudy (anonymous):

then how to get days

OpenStudy (anonymous):

the answer is 578 days.

OpenStudy (anonymous):

no .how the 3 complete the work that much day.it is too long

OpenStudy (anonymous):